OpenStudy (anonymous):

In(x+1)=In(3x+1)-In x

OpenStudy (anonymous):

ln a - ln b =ln (a/b)

OpenStudy (anonymous):

In(x+1)=In(3x+1)-In x In(x+1)=In ((3x+1)/x) hence we can write (x+1)= (3x+1)/x hope u can do further..

OpenStudy (anonymous):

sorry,still cant get it,the ans is